# Yantao Linormal-$\diamondsuit$   Jianbing Zhangnormal-$\diamondsuit$   Zhiyong Wunormal-$\heartsuit$

### 17 Jan 2024 v1

@@ Abstract
Graphical User Interface (GUI) agents are designed to automate complex tasks on digital devices, such as smartphones and desktops. Most existing GUI agents interact with the environment through extracted structured data, which can be notably lengthy (e.g., HTML) and occasionally inaccessible (e.g., on desktops). To alleviate this issue, we propose a visual GUI agent – SeeClick, which only relies on screenshots for task automation. In our preliminary study, we have discovered a key challenge in developing visual GUI agents: GUI grounding – the capacity to accurately locate screen elements based on instructions. To tackle this challenge, we propose to enhance SeeClick with GUI grounding pre-training and devise a method to automate the curation of GUI grounding data. Along with the efforts above, we have also created ScreenSpot, the first realistic GUI grounding dataset that encompasses mobile, desktop, and web environments. After pre-training, SeeClick demonstrates significant improvement in ScreenSpot over various baselines. Moreover, comprehensive evaluations on three widely used benchmarks consistently support our finding that advancements in GUI grounding directly correlate with enhanced performance in downstream GUI agent tasks.  [^1] The model, data and code are available at https://github.com/njucckevin/SeeClick.
@@ 1 Introduction
Developing autonomous agents to assist humans on computing devices has been a persistent goal for artificial intelligence (Shi et al., 2017; Li et al., 2020a; Furuta et al., 2023; Zhou et al., 2023). These Graphical User Interface (GUI) agent systems aim to mimic human interactions in solving complex tasks, thereby enhancing efficiency and reducing manual effort, with examples like Siri and Copilot. Recent advances in Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have significantly propelled the evolution of GUI agents (Gur et al., 2023a; Zhou et al., 2023). These agents interact with the environment through extracted structured texts, e.g., HTML from web pages, then elicit LLM for planning, reasoning, and action gereration (Kim et al., 2023; Zheng et al., 2023).
@@
However, GUI agents depend on structured text face three inherent limitations: (1) Structured text is not always accessible, especially for iOS or desktop applications where acquiring such information is challenging (Shaw et al., 2023); (2) The verbose nature of structured text serves as an inefficient context for LLMs, while also omitting crucial information such as layout, images, and icons (Deng et al., 2023); (3) The variety of structured text - including HTML, DOM, and Android VH - necessitates the curation of task-specific observation and action spaces (Kim et al., 2023; Zhou et al., 2023). These entrenched deficiencies in text-based approaches call for an alternative solution.
@@
In this paper, we propose a visual GUI agent built on Large Vision-Language Models (LVLMs) - SeeClick. Inspired by human interaction with GUIs, SeeClick performs low-level actions like clicking or typing directly by observing interface screenshots. This methodology bypasses the need for interacting with cumbersome structured text, empowering SeeClick as a universal visual agent suitable for various GUI platforms. Building such visual agents involves a foundational challenge: GUI grounding - the capacity to accurately locate screen elements based on instructions, which is absent in current LVLMs. To tackle this challenge, SeeClick enhances LVLM with a GUI grounding pre-training strategy. We devise a method to automate the curation of web grounding data and adapt public mobile UI datasets to obtain mobile grounding data. SeeClick employs the above-curated dataset for continual pre-training of the LVLM, enabling it to accurately locate elements such as text, widgets, and icons in various GUI environments.
@@
Given GUI grounding is a fundamental yet underexplored capacity for GUI agents, we created ScreenSpot, the first realistic GUI grounding evaluation benchmark across various GUI platforms. ScreenSpot contains over 600 screenshots and 1200 instructions from iOS, Android, macOS, Windows, and web environments, and specifically includes both text-based elements and a variety of widgets and icons. Evaluation results confirm SeeClick’s superiority over current LVLMs, validating the effectiveness of GUI grounding pre-training.
@@
Finally, we adapted SeeClick to three downstream agent tasks: MiniWob (Shi et al., 2017), AITW (Rawles et al., 2023), and Mind2Web (Deng et al., 2023). As a purely visual-based agent, SeeClick achieves impressive performance across three tasks. Notably, SeeClick outperforms the visual baseline Pix2Act with only 0.3% training data on MiniWob. Moreover, experimental results on three tasks consistently support our finding that improvement in GUI grounding directly correlates with enhanced agent task performance.
This paper makes the following contributions:
@@
  • We develop a unified visual GUI agent SeeClick, which directly performs clicking and typing actions based on interface screenshots across diverse GUI platforms.
@@
  • We prospectively explore GUI grounding for visual GUI agents, and enhanced SeeClick with our proposed GUI grounding pre-training strategy.
@@ 2.1 Autonomous GUI Navigation
Autonomous GUI navigation requires intelligent agents to tackle complex tasks in various GUIs. Early research investigated training agents by reinforcement learning (Liu et al., 2018; Gur et al., 2018) on the classic web environment MiniWob (Shi et al., 2017). With the recent advancement of LLMs (OpenAI, 2023; Touvron et al., 2023; Xu et al., 2023; Sun et al., 2023; Kim et al., 2023, inter alia). LLM-centric GUI agents have become the dominant paradigm. A line of works focused on prompting closed-source LLMs like ChatGPT and GPT-4, leveraging their powerful reasoning and planning abilities for web tasks, e.g., via in-context learning (Zheng et al., 2023) and self-refine (Kim et al., 2023). Other research explored training LLMs as specialized GUI agents (Gur et al., 2023b; Furuta et al., 2023). Deng et al. (2023) devised a two-stage method for identifying target elements within intricate HTML. Recently, Gur et al. (2023a) designed a method for interacting with websites via synthesizing programs.
@@
However, constrained by the inherent limitation of LLMs to only process text and the complexity of GUIs, converting interfaces into structured text is cumbersome and challenging to standardize across various GUI environments. To address this, several recent efforts have attempted vision-based GUI navigation using GPT-4V (Yan et al., 2023; Gao et al., 2023; Yang et al., 2023). Recent studies explored vision-based UI Agents for web (Shaw et al., 2023) or mobile (Zhan and Zhang, 2023). In this paper, we construct a universal visual GUI agent SeeClick by customizing LVLMs, capable of operating across various GUI platforms.
2.2 Large Vision-Language Models
Recently, researchers have invested tremendous effort in constructing LVLMs capable of jointly processing image and text (Liu et al., 2023a; Zhu et al., 2023; Ye et al., 2023b; Li et al., 2023; Zhang et al., 2023a). LVLMs integrate visual encoding modules, like ViT (Dosovitskiy et al., 2020), with LLMs such as LLaMA (Touvron et al., 2023), inheriting LLMs’ language and reasoning abilities to perform diverse downstream vision-language tasks.
@@ 2.2 Large Vision-Language Models
A series of studies focused on the grounding capabilities of LVLMs (Wang et al., 2023; Bai et al., 2023; Chen et al., 2023a), such as providing bounding boxes for objects while generating responses (Chen et al., 2023b; Peng et al., 2023). Nonetheless, these efforts primarily addressed natural images and did not explore GUI contexts. This paper emphasizes that GUI grounding is pivotal for constructing visual GUI agents and concentrates on exploring LVLMs’ potential in GUI environments.
@@
Figure 1: Overview of our universal visual GUI agent SeeClick. (a) depicts the framework of SeeClick and GUI grounding pre-training. (b) provides examples of ScreenSpot across various GUIs and types of instructions. (c) displays the real-world application of SeeClick on downstream web agent task.
2.3 Mobile UI Understanding
As a vital aspect of daily life, mobile UI understanding has garnered widespread attention from researchers. One category of work explored multimodal tasks within UI contexts (Li et al., 2021; Bai et al., 2021; He et al., 2021), such as widget captioning (Li et al., 2020b), screen summarization (Wang et al., 2021), and UI object detection (Chen et al., 2020). Several studies explored grounding elements in mobile UI based on instructions (Li et al., 2020a; Burns et al., 2022; Li et al., 2021; Li and Li, 2022), serving as a foundation for mobile UI automation. Most mentioned studies rely on Android’s View Hierarchy to represent interfaces, while (Li and Li, 2022) use Regions of Interest and (Zhang et al., 2023b) use screenshots. Moreover, most efforts employed RICO (Deka et al., 2017) for research, an Android apps dataset collected in 2017. This work introduces an up-to-date, realistic grounding benchmark encompassing various GUI platforms, and underscores grounding as a foundation for visual GUI agents.
@@ 3 Approach
In our preliminary study, we have discovered a key challenge in developing visual GUI agents: GUI grounding, the capacity to locate screen elements based on instructions. Recent advancements in LVLMs have claimed visual grounding capability on natural images (Chen et al., 2023b; Peng et al., 2023; Wang et al., 2023), which is able to point out objects based on natural language description. However, GUI screenshots significantly differ from natural images, typically containing dense text and numerous icons and widgets. These differences impair existing LVLMs’ grounding performance in GUI contexts (Bai et al., 2023; Chen et al., 2023a), limiting their potential as visual GUI agents.
This paper seeks to endow LVLMs with GUI grounding proficiency, aiming to construct a universal visual GUI agent capable of executing instructions that only rely on screenshots. The overview of the proposed LVLM SeeClick is presented in Figure 1(a). Next, we introduce the birth of SeeClick, including the formalization of GUI grounding task, the construction of continual pre-training data, and training details.
@@ 3.1 GUI grounding for LVLMs
As GUI grounding is the core capability of SeeClick, we first elucidate how to train LVLM for language generation to perform grounding tasks. Given an interface screenshot $s$ and a collection of its elements $\{(x_{i},y_{i})|_{i}\}$, where $x_{i}$ denotes the textual description of the $i$-th element and $y_{i}$ indicates the element’s location (represented as a bounding box or point). As depicted in Figure 1(a), LVLM predicts the location of the element $y$ based on the interface screenshot $s$ and its textual description $x$, i.e. calculating $p(y|s,x)$.
A potential challenge is how LVLMs predict numerical coordinates in a language generation format. Traditional methods (Chen et al., 2021; Wang et al., 2023; Shaw et al., 2023) divided the image into 1000 bins, and creating a new 1,000-token vocabulary $\{<p0>,<p1>,...,<p999>\}$ to represent the x and y coordinates. In this work, we adopt a more intuitive manner used in LVLMs (Chen et al., 2023b; Bai et al., 2023), treating numerical values as natural language without any additional tokenization or pre-/post-processing. For instance, in Figure 1(a), for a smartphone screenshot and the instruction “View the new album of Jony J”, we craft a user query prompt: “In the UI, where should I click if I want to <instruction>?”. Subsequently, we normally compute the cross-entropy loss between the model output and the groundtruth “click (0.49, 0.40)” to optimize the LVLM.
@@
3.2 Data Construction
We primarily train SeeClick using three collections of data: web UI data crawled from the internet, mobile UI data reorganized from public datasets, and general vision-language instruction-following data.
Web Data. Web UIs, featuring a variety of layouts and design styles across websites, are ideal for training LVLMs to learn universal recognition, understanding, and grounding capabilities across different GUIs. We collected approximately 300k web pages from the latest Common Crawl repository  [^2] https://commoncrawl.org/ to serve as our training data for web UI. For each webpage, we collect two types of elements from the HTML code: (1) elements that display visible text content; and (2) elements with a special “title” attribute that display descriptive text when hovering. This method ensures that we gather a series of interactable elements and their corresponding text instructions, while encompassing a wide range of texts and icons. Figure 2 provides an example of a web page, its associated simplified HTML code and extracted two type elements. Next, we treat the text or “title” content of each element as the instruction, constructing the GUI grounding data with screenshot $s$, instruction $x$, and the element’s location $y$. We also include an inverse task, similar to web OCR, where given an interface screenshot $s$ and the location $y$ of an element, the model predicts the text description $x$ of the element.
Mobile Data. For mobile UI, we include three types of data: widget captioning, mobile UI grounding, and mobile UI summarization. Widget Captioning aims to automatically generate language descriptions for mobile UI elements; for example, producing the description “play music” for the play button on a music player interface. We utilize the training split of the dataset provided in (Li et al., 2020b), containing nearly 20k screenshots, 40k widgets and 100k descriptions. We derive mobile UI grounding data by inverting the widget captioning task, treating language descriptions as instructions and predicting widget locations from screenshots and instructions. To improve element diversity, we also incorporate the automatically generated elements and instructions from RICO (Li et al., 2020a). The mobile UI grounding data involve diverse elements and instructions, aids SeeClick in generalizing grounding capabilities to various GUI scenarios. We finally include mobile UI summarization data (Wang et al., 2021) to enhance overall interface comprehension.
General Data. To maintain LVLM’s general capacities on natural images and its planning and reasoning abilities for downstream agent tasks, we incorporate general vision-language data for training SeeClick. We utilize the 158k instruction-following samples collected with GPT-4 from LLaVA (Liu et al., 2023a), covering conversation, detailed description, and complex reasoning.
@@ 3.3 Training Details
We built SeeClick through continual pre-training on a recent advanced LVLM, Qwen-VL (Bai et al., 2023). Qwen-VL inherently possesses grounding capabilities and a higher resolution (448*448), which has been proven beneficial for understanding text-rich images (Lee et al., 2023; Ye et al., 2023a). We train Qwen-VL on the constructed dataset (Section 3.2) for about 10k steps (around 1 epoch) to obtain our GUI base model SeeClick. During training, we employ LoRA (Hu et al., 2021) to fine-tune both the visual encoder and LLM. We adopt AdamW as the optimizer and use a cosine annealing scheduler with an init learning rate of 3e-5 and a global batch size of 64. All training takes around 24 hours on 8 NVIDIA A100 GPUs.
Table 1: Results of different LVLMs on ScreenSpot. The best results in each column are bold. Benefit from efficient GUI grounding pre-training, SeeClick significantly enhanced LVLMs’ ability to locate GUI elements following instructions, and surpassed the strong baseline CogAgent with a smaller model size.
@@
MiniGPT-v2 7B 7.6% 5.2% 4.5% 2.9% 5.5% 0.0% 4.3%
Qwen-VL 9.6B 9.4% 4.3% 6.4% 5.3% 0.0% 0.0% 4.2%
Fuyu 8B 40.6% 1.6% 33.6% 6.7% 48.4% 2.9% 22.3%
CogAgent 18B 66.5% 26.7% 73.7% 19.3% 78.0% 21.4% 47.6%
SeeClick 9.6B 75.9% 48.7% 72.7% 26.4% 69.2% 21.4% 52.4%
4 GUI Grounding Benchmark: ScreenSpot
We recognize GUI grounding proficiency as essential for constructing a universal visual GUI agent. However, due to the constrained capabilities of earlier vision-language models, previous research has paid limited attention to GUI grounding, with the few existing studies (Li et al., 2021; Li and Li, 2022; Zhang et al., 2023b) largely confined to Android environments.
To address this research gap, we introduce ScreenSpot, an up-to-date, realistic grounding evaluation benchmark encompassing various GUI platforms. It is designed to assess vision-language models’ ability to locate elements based on human instructions (Figure 1(b) provides some examples). ScreenSpot has two distinctive features: (1) Various GUI platforms. It includes over 600 interface screenshots from mobile (iOS, Android), desktop (macOS, Windows), and web platforms, along with more than 1200 instructions and corresponding actionable elements (as shown in Figure 3 left); (2) Icons/Widgets. Considering that locating icons or widgets might be more challenging than text elements on GUIs, we include a substantial number of icon/widget type elements in each GUI (as depicted in Figure 3 right). See Appendix A for more examples.
To evaluate vision-language models’ effectiveness in real-world scenarios, ScreenSpot is carefully curated to ensure the samples in ScreenSpot are novel and not included in existing training resources. We recruited experienced annotators to collect new GUI interfaces and label instructions along with the bounding boxes for actionable elements. For mobile and desktop, annotators were asked to select commonly used apps and operations; for web pages, we chose several types of websites (development, shopping, forum, and tools) from the recent web environment WebArena (Zhou et al., 2023) for annotation.

### 23 Feb 2024 v2

@@ Abstract
Graphical User Interface (GUI) agents are designed to automate complex tasks on digital devices, such as smartphones and desktops. Most existing GUI agents interact with the environment through extracted structured data, which can be notably lengthy (e.g., HTML) and occasionally inaccessible (e.g., on desktops). To alleviate this issue, we propose a novel visual GUI agent – SeeClick, which only relies on screenshots for task automation. In our preliminary study, we have discovered a key challenge in developing visual GUI agents: GUI grounding – the capacity to accurately locate screen elements based on instructions. To tackle this challenge, we propose to enhance SeeClick with GUI grounding pre-training and devise a method to automate the curation of GUI grounding data. Along with the efforts above, we have also created ScreenSpot, the first realistic GUI grounding benchmark that encompasses mobile, desktop, and web environments. After pre-training, SeeClick demonstrates significant improvement in ScreenSpot over various baselines. Moreover, comprehensive evaluations on three widely used benchmarks consistently support our finding that advancements in GUI grounding directly correlate with enhanced performance in downstream GUI agent tasks.  [^1] The model, data and code are available at https://github.com/njucckevin/SeeClick.
@@ 1 Introduction
A perennial topic in machine intelligence is the development of Graphical User Interface (GUI) agent systems, like Siri and Copilot, to automate complex tasks on computing devices, thereby reducing human workload (Shi et al., 2017; Li et al., 2020a). Recent advances in Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023) have significantly propelled the evolution of GUI agents (Gur et al., 2023; Zhou et al., 2023). These agents interact with the environment by interpreting structured texts, e.g., HTML from webpages, then elicit LLM for planning, reasoning, and execution (Kim et al., 2023; Zheng et al., 2023).
Figure 1: Text-based agents select target elements from structured texts, occasionally augmented with screenshots. SeeClick employs a vision-based methodology to predict action locations solely relying on screenshots.
@@
However, GUI agents depend on structured text face three inherent limitations: (1) Structured text is not always accessible, especially for iOS or desktop applications where acquiring such information is challenging (Shaw et al., 2023); (2) The verbose nature of structured text constitutes an inefficient context for LLMs, while also omitting crucial information such as layout, images, and icons (Deng et al., 2023); (3) The variety of structured text - including HTML, DOM, and Android VH - necessitates the curation of task-specific observation and action spaces (Kim et al., 2023; Zhou et al., 2023). These entrenched deficiencies in text-based approaches call for an alternative solution.
@@
In this paper, we introduce SeeClick, a visual GUI agent built on Large Vision-Language Models (LVLMs). Inspired by human interaction with GUIs, as illustrated in Figure 1, SeeClick is designed to perform low-level actions like clicking or typing directly by observing interface screenshots. This innovative approach bypasses the interaction with cumbersome structured text, empowering SeeClick to universally adapt to various GUI platforms. Building such visual agents presents a foundational challenge: GUI grounding - the capacity to accurately locate screen elements based on instructions, which is absent in current LVLMs.To tackle this challenge, SeeClick enhances LVLM with a GUI grounding pre-training strategy. We devise a method to automate the curation of web grounding data and adapt public mobile UI datasets to obtain mobile grounding data. SeeClick employs the above-curated dataset for continual pre-training of the LVLM, enabling it to accurately locate elements such as text, widgets, and icons in various GUI environments.
@@
Given GUI grounding is a fundamental yet underexplored capacity for GUI agents, we establish ScreenSpot, the first realistic GUI grounding evaluation benchmark across various GUI platforms. ScreenSpot contains over 600 screenshots and 1200 instructions from iOS, Android, macOS, Windows, and webpages, and specifically includes both text-based elements and a variety of widgets and icons. Evaluation results confirm SeeClick’s superiority over current LVLMs, validating the effectiveness of GUI grounding pre-training.
@@
Finally, we adapt SeeClick to mobile and web agent tasks, including MiniWob (Shi et al., 2017), AITW (Rawles et al., 2023), and Mind2Web (Deng et al., 2023). As a purely vision-based agent, SeeClick achieves impressive performance. It surpasses the strong visual baseline Pix2Act while utilizing merely 0.3% training data. Moreover, experimental results on these three benchmarks consistently support our findings that improvement in GUI grounding directly correlates with enhanced agent task performance.
Our main contributions are as follows:
@@
  • We develop a unified visual GUI agent SeeClick, which solely relies on interface screenshots to perform clicking and typing actions across diverse GUI platforms.
@@
  • We prospectively explore GUI grounding for visual GUI agents, and enhanced SeeClick with proposed GUI grounding pre-training strategy.
@@ 2 Related work
Autonomous GUI Navigation Early research explored task automation in simplified web (Shi et al., 2017; Liu et al., 2018; Gur et al., 2018) and mobile UI (Li et al., 2020a; Burns et al., 2022; Li and Li, 2022). With LLM advancements (OpenAI, 2023; Touvron et al., 2023; Xu et al., 2023; Sun et al., 2023; Wu et al., 2024, inter alia), LLM-centric agents have become the dominant paradigm. A line of works focused on prompting ChatGPT and GPT-4 for web tasks, via in-context learning (Zheng et al., 2023) and self-refine (Kim et al., 2023). Other research explored training LLMs as specialized agents. Deng et al. (2023) devised a two-stage method for identifying target elements within intricate HTML. Gur et al. (2023) proposed to interact with websites via programming.
@@
Given the constraints of LLM to only process text, recent efforts have attempted vision-based GUI navigation (Shaw et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023). These methods primarily utilize GPT-4V (Yan et al., 2023; Gao et al., 2023) and also require GUI metadata as input (Yang et al., 2023a; Zheng et al., 2024). In this paper, we construct a universal visual GUI agent SeeClick by customizing open-sourced LVLM, capable of operating across various GUI platforms without needing any GUI metadata.
@@
Large Vision-Language Models Recent research has invested tremendous effort in constructing LVLMs capable of jointly processing image and text (Liu et al., 2023a; Zhu et al., 2023; Ye et al., 2023; Li et al., 2023), integrating vision encoders with LLMs through connecting layers, inheriting LLMs’ linguistic and reasoning skills to perform vision-language tasks. A series of studies focused on grounding with LVLMs (Wang et al., 2023; Bai et al., 2023; Chen et al., 2023a), such as providing bounding boxes for objects when generating responses (Chen et al., 2023b; Peng et al., 2023). Nonetheless, these efforts primarily addressed natural images and did not explore GUI contexts. This paper focuses on grounding in GUIs and explores the potential of LVLMs as visual agents.
@@
Figure 2: Overview of our universal visual GUI agent SeeClick. (a) depicts the framework of SeeClick and GUI grounding pre-training. (b) provides examples of ScreenSpot across various GUIs and types of instructions. (c) displays the real-world application of SeeClick when adapted to downstream web agent tasks.
@@ 3 Approach
Our preliminary study highlights a major challenge in developing visual GUI agents: GUI grounding, the capacity to locate screen elements based on instructions. Although recent LVLMs have claimed grounding capability on natural images (Bai et al., 2023; Wang et al., 2023), GUI screenshots differ significantly with dense text and numerous icons and widgets. These differences impair existing LVLMs’ grounding performance in GUI contexts and limit their potential as visual GUI agents.
This paper seeks to harness LVLMs with GUI grounding skills, paving the way for a visual GUI agent that executes instructions only relying on screenshots. As presented in Figure 2, SeeClick is a foundational model for GUIs, and tailored for adaption to agent tasks. Next, we introduce the birth of SeeClick, including the formalization of GUI grounding task, the construction of continual pre-training data, and training details.
@@ 3.1 GUI grounding for LVLMs
As GUI grounding is the core capability of SeeClick, we first elucidate how to train LVLM for language generation to perform grounding tasks. Given an interface screenshot $s$ and a collection of elements $\{(x_{i},y_{i})|_{i}\}$ on it, where $x_{i}$ denotes the textual description of the $i$-th element and $y_{i}$ indicates the element’s location (represented as a bounding box or point). As depicted in Figure 2(a), LVLM predicts the location of the element $y$ based on the interface screenshot $s$ and its textual description $x$, i.e. calculating $p(y|s,x)$.
A potential challenge is how LVLMs predict numerical coordinates in a language generation format. Previous studies (Chen et al., 2021; Wang et al., 2023; Shaw et al., 2023) divide the image into 1000 bins, and creating a new 1,000-token vocabulary $\{<p0>,<p1>,...,<p999>\}$ to represent the x and y coordinates. In this work, we adopt a more intuitive manner used in LVLMs (Chen et al., 2023b; Bai et al., 2023), treating numerical values as natural languages without any additional tokenization or pre-/post-processing. For instance, in Figure 2(a), for a smartphone screenshot and the instruction “View the new album of Jony J”, we craft a query prompt: “In the UI, where should I click if I want to <instruction>?”. Subsequently, we normally compute the cross-entropy loss between the model output and the ground truth “click (0.49, 0.40)” to optimize the LVLM.
3.2 Data Construction
We train SeeClick using three collections of data: web UI data crawled from the internet, mobile UI data reorganized from public datasets and general vision-language instruction-following data.
Web Data. Web UIs, featuring a variety of layouts and design styles across websites, are ideal for training LVLMs’ general recognition and grounding capabilities across different GUI contexts. We collect approximately 300k web pages from the latest Common Crawl repository to serve as our training data for web UI. For each webpage $s$, we collect two types of elements from the HTML code as exemplified in Figure 3: (1) elements that display visible text content; and (2) elements with a special “title” attribute that display descriptive text when hovering. This method ensures that we gather a series of interactable elements $y$ and their corresponding instructions $x$, while encompassing a wide range of text and icon elements. In addition to the grounding task $p(y|s,x)$, we also include web OCR task $p(x|s,y)$, predicting text description based on coordinates.
@@ 3.2 Data Construction
Mobile Data. For mobile UI, we include three types of data: widget captioning, mobile UI grounding, and mobile UI summarization. The widget captioning dataset provides language descriptions for mobile UI elements; for example, the description “play music” for the play button on a music player interface. We utilize the training split of the dataset provided by (Li et al., 2020b), containing nearly 20k screenshots, 40k widgets, and 100k descriptions. We derive mobile UI grounding data by reversing the process of widget captioning, treating language descriptions as instructions and corresponding widgets as target elements. To improve diversity, we also incorporate the automatically collected elements and instructions from RICO (Li et al., 2020a). The mobile data involves diverse elements and instructions, facilitating the generalization of SeeClick’s grounding proficiency to diverse GUI contexts. We finally include mobile UI summarization data (Wang et al., 2021) to enhance overall interface comprehension.
General Data. To maintain LVLM’s general capacities on natural images, we incorporate the general vision-language instruction-following data from LLaVA (Liu et al., 2023a), covering conversation, detailed description, and complex reasoning.
@@ 3.3 Training Details
We build SeeClick through continual pre-training on a recent advanced LVLM, Qwen-VL (Bai et al., 2023), which possesses grounding capabilities and a higher resolution of 448*448. We train Qwen-VL on the dataset we constructed (as described in Section 3.2) for about 10k steps (around 1 epoch) to obtain our GUI base model SeeClick. During training, we employ LoRA (Hu et al., 2021) to fine-tune both the visual encoder and LLM. Further details and task examples are provided in Appendix A.
@@
MiniGPT-v2 7B 8.4% 6.6% 6.2% 2.9% 6.5% 3.4% 5.7%
Qwen-VL 9.6B 9.5% 4.8% 5.7% 5.0% 3.5% 2.4% 5.2%
GPT-4V - 22.6% 24.5% 20.2% 11.8% 9.2% 8.8% 16.2%
Fuyu 8B 41.0% 1.3% 33.0% 3.6% 33.9% 4.4% 19.5%
CogAgent 18B 67.0% 24.0% 74.2% 20.0% 70.4% 28.6% 47.4%
SeeClick 9.6B 78.0% 52.0% 72.2% 30.0% 55.7% 32.5% 53.4%
Table 1: Results of different LVLMs on ScreenSpot. The best results in each column are highlighted in bold. Benefiting from efficient GUI grounding pre-training, SeeClick significantly enhanced LVLMs’ ability to locate GUI elements following instructions, and surpassed the strong baseline CogAgent with a smaller model size.
4 ScreenSpot: A Grounding Benchmark
We recognize GUI grounding proficiency as essential for constructing visual GUI agents. However, the constrained capabilities of earlier vision-language models resulted in limited attention, with scant research (Li et al., 2021; Li and Li, 2022; Zhang et al., 2023) largely confined to an Android dataset (Deka et al., 2017) collected in 2017.
To address this research gap, we introduce ScreenSpot, an up-to-date, realistic grounding evaluation benchmark encompassing various GUI platforms. It is designed to assess vision-language models’ ability to locate screen elements based on instructions (Figure 2(b) provides some examples). ScreenSpot has two distinctive features: (1) Various GUI platforms. It includes over 600 interface screenshots from mobile (iOS, Android), desktop (macOS, Windows), and web platforms, along with 1200+ instructions and corresponding actionable elements; (2) Icons/Widgets. ScreenSpot includes a substantial number of icons and widgets in each GUI, which is more challenging to locate than texts (statistics are in Figure 4). See Appendix B for annotation details and examples.
To measure models’ effectiveness in real-world scenarios, ScreenSpot is carefully curated to ensure the samples are novel and not included in existing training resources. We recruited experienced annotators to collect GUI interfaces and label instructions along with the bounding boxes for actionable elements. For mobile and desktop, annotators were asked to select commonly used apps and operations; for web, we chose several types of websites (development, shopping, forum, and tools) from the web environment WebArena (Zhou et al., 2023).